# On the extension of a class of Hermite bivariate interpolation problems

### v1

@@ Abstract
We characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space. This result extends an earlier theorem (2000) by one of the authors concerning the case . The latter theorem, in turn, can be regarded as a natural generalization of a classical theorem of Severi (1921).
Key words: Fundamental polynomial, algebraic curve, -correct set, -solvable set.
@@ 1 Introduction
In this paper, we characterize the sets of solvability for Hermite bivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1). This extends a previous result (Theorem 1.20) that addresses the case of . The latter theorem can, in turn, be regarded as a natural generalization of a classical theorem of Severi (Theorem 1.21).
One should mention that the correctness of Hermite problems was considered first in algebraic geometry by M. Nagata [6]. In approximation theory the investigation was initiated by G. G. Lorentz and R. A. Lorentz [4].
@@
1.1 Lagrange bivariate interpolation
Denote the space of bivariate polynomials of total degree with real or complex coefficients:
Let us fix a set of distinct nodes in the plane:
The problem of finding a polynomial satisfying the conditions
for a given data is called Lagrange interpolation problem.
Definition 1.1.
A set of nodes is called -solvable if for any data there exists a polynomial (not necessarily unique) , satisfying the conditions (1.2).
Denote by the restriction of on the set
A polynomial is called an -fundamental polynomial for if
We denote this -fundamental polynomial by
Let us mention that sometimes we call -fundamental a polynomial satisfying the conditions
since it equals a non-zero constant times
Definition 1.2.
A set of nodes is called -independent if each node has an -fundamental polynomial. Otherwise, it is called -dependent.
If a set is -independent then the following Lagrange formula gives a solution to the problem (1.2):
Thus we obtain:
Proposition 1.3.
A node set is -independent if and only if it is -solvable.
Fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, is
Definition 1.4.
A set of nodes is called -correct if for any data there exists a unique polynomial , satisfying the conditions (1.2).
A necessary condition of -correctness is:
For the sets of nodes satisfying the above condition we have
Proposition 1.5.
Assume that Then the following conditions are equivalent:
The set is -correct.
The set is -solvable.
A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation . In particular, by we denote a linear polynomial from and the line defined by the equation Similarly by we denote a quadratic polynomial from and the conic defined by the equation
1.2 Hermite bivariate interpolation
Hermite interpolation problem is determined by a pair where is a set of nodes given by (1.1) and is a scheme, i.e., a sequence of multiplicities of the nodes associated with the degree
@@ 1.2 Hermite bivariate interpolation
Denote
In the problem the interpolation parameters at a given node are the value of a polynomial and its partial derivatives up to the order
@@
is called Hermite interpolation problem.
We express briefly the conditions (1.4) as follows
Nodes whose multiplicity is equal to one are called simple. At these nodes only the value of a polynomial is interpolated in the conditions (1.4).
It is sometimes convenient to allow nodes of multiplicity in the problem In fact, such a node can be ignored because there is no interpolation condition.
@@
A problem is called -independent if for each parameter in (1.4) there is an -fundamental polynomial.
@@
Denote
We have that the number of interpolation conditions at a node in (1.4) equals to Hence the total number of the conditions equals to
It is easily seen that fundamental polynomials are linearly independent. Also we have that Therefore a necessary condition of -independence, or -solvability, of the problem is
@@
A scheme satisfying the above condition we call an exact scheme.
Below we have the analog of Proposition 1.5 for the Hermite interpolation.
@@
Assume that a scheme is exact. Then the following conditions are equivalent:
The problem is -correct.
@@
Indeed, note that is the number of interpolation conditions in (1.4) and each condition can be considered as a linear equation with respect to unknown coefficients of Thus we get a linear homogeneous system, where the number of equations is less than the number of unknowns. As we know such systems have non-trivial solutions.
In the simplest case the above proposition implies
Lemma 1.12.
For any five points, there exists a conic that passes through all of them.
For a -scheme we define the exact scheme
@@
The direction “only if” is evident. For the direction “if”, in view of Proposition 1.10, it is enough to prove that there exists a (simple) node such that the problem where and is -solvable. For this, in view of -solvability of we need just an -fundamental polynomial of the simple node
It is easily seen that the polynomial is the desired -fundamental polynomial of ∎
@@
Let be a problem, where is given by (1.1), and be a curve of degree Then we define the scheme as follows
@@
Denote by the number of nodes from on the curve counted with the multiplicities from
Note that in what follows, this notation, as well as the previous definition, mainly will be used in cases where the curve is a line or a conic.
We will need the following well-known proposition (see e.g. [2], lemma 4):
@@
Let be a scheme with and be a set of nodes with where is a line. Then
@@
The following proposition we will use for lines and (irreducible) conics.
@@
Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then for any curve of degree we have that
@@
Let the problem be -solvable for a set of nodes Assume, by way of contradiction, that for a curve of degree we have that
@@
Consider the following scheme
@@
Then let us check that is -scheme. Indeed, by using (1.7) and (1.9), we get
@@
Now one readily obtains for the polynomial that
This, in view of Proposition 1.10, contradicts the -correctness of the problem ∎
@@
Let be a scheme with non-increasing sequence of multiplicities. If the problem is -solvable for a set of nodes then:
@@
Note, that by taking into account the above condition (i) for the lines passing through any two nodes, we obtain
@@
Let be a scheme with non-increasing sequence of multiplicities. If the problem is solvable for a set of nodes then
Now we are in a position to present the result (slightly reformulated), which we will extend.
Theorem 1.20 ([2], th. 9).
@@
Then the problem is -solvable if and only if there is no line passing through more than nodes of
This theorem can be regarded as a natural generalization of the following classical theorem of Severi (1921).
Theorem 1.21 (Severi [7], th. 9).
Let be a scheme. The problem is solvable for any node set if and only if
@@
For any scheme with non-increasing sequence of multiplicities, we have that
Indeed, consider the problem where in no three nodes are collinear. Then according to Theorem 1.20 the problem is -solvable. Hence is -scheme.
@@
Suppose is a set of nodes with where is an irreducible conic. Then we have that
@@
The above inequality, in view of irreducibility of implies that the -condition is satisfied. Thus, in view of Theorem (1.20), we obtain that the problem is -solvable. Similarly we obtain that the problem is -solvable.
Now consider the following linear spaces:
Note that In view of … we have that
@@
Then we readily get that implies that While implies that
Both these schemes are exact. According to the above proof the problems are also solvable, provided with any irreducible conic. The latter condition in this case means that no three nodes are collinear. Thus the problems are correct:
This holds also with the scheme replaced by Therefore Proposition 1.23 is valid also in the case since the statement reduces to
@@
Note that according to Corollary 1.22 for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
In the main result of this paper below we consider instead the condition which, in combination with does not imply that is -scheme. Indeed, for a counterexample, we just add a fifth multiplicity to the schemes and of the above remark.
This is why in the next theorem, unlike Theorem 1.20, we have the condition: is -scheme.
2 The result
@@
Then the problem is -solvable if and only if
there is no line passing through more than nodes of ,
there is no conic passing through more than nodes of
@@
The direction “only if” follows from Corollary 1.18. Let us prove the direction “if”.
Note that the above condition (ii) means that there is no conic passing through all the nodes of
@@
In view of Corollary 1.19 we have that
@@
Indeed, the only - scheme, in the case is the Taylor scheme which does not satisfy (2.2). While if then the conditions (2.2) and (2.4) imply that which is not - scheme.
@@
In the case there is only one - scheme satisfying the conditions (2.2) and (2.5). According to the condition (ii) of Theorem, the six simple nodes of the set do not belong to a conic. Since is exact scheme, by Proposition 1.5 (iii), we get that is -correct.
@@
Step 2. Assume that there is a line which passes through nodes of i.e., the following condition is satisfied:
@@
Now let us check that is - scheme. Indeed, by using (2.7), we get
The condition (2.1) for readily follows from the fact that passes through at least two nodes of
Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing nodes of Assume conversely that there is such a line Then evidently and, in view of the relation
we obtain that contains nodes from Since contains nodes both from and we conclude that the line does not intersect at a node of Therefore, the (reducible) conic passes through nodes of which contradicts the condition (ii) of Theorem.
Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing nodes of In view of (2.9) this means that passes through all the nodes with
Note that if passes through at least three nodes of then in (2.9) we would obtain and hence the -condition is satisfied. It remains to consider the case where passes through exactly two nodes, say, In this case, in view of (2.6), we have that According to (2.5) we get that Hence we have that Thus passes through all the nodes of which is contradiction.
@@
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
@@
and the nodes of correspond to the last ones in
Here, by making use of Remark 1.14, we can also assume that the set contains no node belonging to the line
@@
Let us check first that is exact scheme. Indeed, in view of (2.7), we have that
Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.10, (iii), that where and prove that
@@
Now, Proposition 1.16 implies that where
Now, in view of (2.10) and Proposition 1.10, we get and hence
@@
Step 3. Assume that there is a line passing through at least nodes of Consider the following scheme (see (1.6))
@@
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
@@
Again, by making use of Rremark 1.14, we can assume also that the set contains no node belonging to the line
Then denote by a set of simple nodes in
@@
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
@@
Now, in view of (2.13) and Proposition 1.10, we get and hence
@@
From now on, in view of Step 3, we may assume that there are no three collinear nodes in the set
Step 4. First note that the condition (ii) of Theorem, in view of Lemma 1.12, implies that
Consider a conic passing through the nodes Since no three nodes are collinear the conic is irreducible.
@@
Now, assume conversely, that the inequality holds for some
Then, in view of (2.3), we have that and therefore
@@
Thus, according to Theorem 1.20, the problem is -solvable.
Now, Proposition 1.13 implies that there is a set consisting of simple nodes, such that the problem
@@
By making use of Remark 1.14 we can assume that the set contains no node belonging to the conic
Next, denote by a set of simple nodes in
@@
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.10, (iii), that where and prove that
@@
Now, in view of (2.15) and Proposition 1.10, we get and hence

### v2

# On the extension of a class of Hermite multivariate interpolation problems

@@ Abstract
We characterize the sets of solvability for Hermite multivariate interpolation problems when the sum of multiplicities is at most , with the degree of the polynomial space. This result extends an earlier theorem (2000) by one of the authors concerning the case . The latter theorem, in turn, can be regarded as a natural extension of a classical Theorem of Severi (1921).
Key words: Hermite interpolation, fundamental polynomial, algebraic curve, -correct set, -solvable set.
@@ 1 Introduction
In this paper, we characterize the sets of solvability for Hermite multivariate interpolation problems when the sum of multiplicities is at most , where denotes the degree of the polynomial space (see Theorem 2.1). This result extends a previous one (Theorem 1.18) that addresses the case of . The latter theorem can, in turn, be regarded as a natural extension of a classical Theorem of Severi (Theorem 1.19).
One should mention that the correctness of Hermite problems was considered first in algebraic geometry by M. Nagata [7]. In approximation theory the investigation was initiated by G. G. Lorentz and R. A. Lorentz [4], [5].
@@
1.1 Hermite multivariate interpolation
Hermite interpolation by multivariate algebraic polynomials is considered. The interpolation parameters are the values of a function and its partial derivatives up to some orders at given points.
Denote for and the multi-index
Set for convenience
The space of polynomials in variables of total degree is denoted by
The coordinates of points and coefficients of polynomials can be taken either real or complex, thus leading to a real or complex interpolation problem, respectively.
Consider a set of distinct interpolation points
Consider also a scheme, i.e., a sequence of multiplicities of the points associated with the degree of interpolating polynomials
@@ 1.1 Hermite multivariate interpolation
In Hermite interpolation problem where is the dimension, the interpolation parameters at a given point are the value of a polynomial and its partial derivatives up to the order
@@
is called Hermite interpolation problem
From now on we express briefly the conditions (1.4) as follows
Points whose multiplicity is equal to one are called simple. At these points only the value of a polynomial is interpolated in the conditions (1.4).
It is sometimes convenient, for example in proofs, to allow points of multiplicity in the problem In fact, such a point can be ignored because there is no interpolation condition.
@@
Let us mention that sometimes we call -fundamental a polynomial, which is a solution to the problem (1.4) where since it equals a non-zero constant times
@@
A problem is called -independent if for each parameter in (1.4) there exists an -fundamental polynomial.
@@
We have that the number of interpolation conditions at a point in (1.4) equals to (see 1.1). Hence the total number of the conditions equals to
It is easily seen that fundamental polynomials are linearly independent. Therefore a necessary condition of -independence, or -solvability, of the problem is
@@
A scheme satisfying the above condition we call exact scheme.
@@
Assume that a scheme is exact. Then the -correctness of the problem is equivalent to each of the following conditions:
@@
Proof.
Indeed, each of the conditions in (1.4) can be considered as a linear equation with respect to unknown coefficients of Thus for exact scheme the conditions (1.4) reduce to a square linear system (see (1.6)). Now, the condition (i) means that the linear system is consistent for any right-hand side values, while the condition (ii) means that the homogeneous linear system has a single trivial solution. It is well known that in the case of a square system each of these properties implies that the system has a unique solution for any right-hand side values. The later property means that the problem is -correct. ∎
@@
Proof.
Indeed, in this case the conditions (1.4) reduce to a linear homogeneous system, where the number of equations is less than the number of unknowns. As we know such systems have non-trivial solutions. ∎
Thus, for the -scheme we obtain
Corollary 1.7.
For any five points in the plane, there exists a conic that passes through all of them.
For a -scheme we define the associated exact scheme
@@
The direction “only if” is evident. For the direction “if”, in view of Proposition 1.5(i), it is enough to prove that there exists a (simple) point such that the problem where and is -solvable. For this, in view of the -solvability of we need just to have an -fundamental polynomial of the simple point
Then, it is easily seen that the polynomial is the desired -fundamental polynomial of ∎
Denote
Corollary 1.9.
Suppose that the problem is -solvable. Then we have that
Proof.
Remark 1.10 ([3], rem.1.14).
@@
If the problem is -solvable, with all belonging to an -dimensional linear subspace then so is the problem where the projection of on is
Until the end of this subsection, we study the bivariate case.
In this case we use the following simplified notation:
In this case instead of (1.1) it is more convenient to use the notation
Then the equality (1.5) and the second equality of (1.2) become:
A plane algebraic curve is the zero set of some bivariate polynomial of degree To simplify notation, we shall use the same letter, say , to denote the polynomial and the curve given by the equation . In particular, by we denote a linear polynomial from and the line defined by the equation Similarly, by we denote a quadratic polynomial from and the conic defined by the equation
@@
Let be a problem, where is a set given by (1.3), and be a plane curve of degree Then we define the scheme as follows
@@
Denote by the number of points from on the curve counted with the multiplicities from
Note that this notation, as well as the previous definition, mainly will be used in the cases where the curve either is a line or a conic.
For denote
We will need the following well-known proposition:
@@
Let be a scheme with and be a set of points with where is a line. Then
where and
Corollary 1.13.
Suppose is a problem, where is -scheme. Suppose also is a line with Then the problem is -solvable if and only if the problem is -solvable, where
1.2 Some results in the multivariate case
@@
The following proposition will be used for lines and conics.
@@
Let be a scheme. If the problem is -solvable for a set of points then for any plane curve of degree we have that
@@
First consider the case Let the problem be -solvable for a set of points Assume, by way of contradiction, that for a curve of degree we have that
@@
Now, consider the following scheme
@@
Then let us check that is -scheme. Indeed, by using (1.10) and (1.12), we obtain
@@
Now we readily get that the polynomial satisfies the condition:
Now consider the case of general dimension: Assume that does not satisfy the condition, i.e., there is a plane curve passing through more than of its points. Consider the respective subproblem restricted to the interpolation points belonging to the curve As we proved above, this subproblem is not -solvable in the case of Therefore the original problem also is not -solvable. ∎
@@
Let be a scheme. If the problem is -solvable for a set of points then:
@@
Definition 1.16.
A scheme is called regular if there exists a set of points such that the problem is -independent.
By taking into account the above condition (i) for the lines passing through any two points, and the condition (ii) for the conics passing through any five points, one readily obtains
@@
Let be a scheme with non-increasing sequence of multiplicities. If is regular then we have that
Now we are in a position to present the result, which we are going to extend. Actually, it is slightly reformulated (see [2], 2nd par. after the proof of Thm.9).
Theorem 1.18 ([2], Thm.9).
@@
Then the problem is -solvable if and only if there is no line passing through more than points of
This theorem can be regarded as a natural extension of the following classical Theorem of Severi (1921).
Theorem 1.19 (Severi [8], Thm.9).
Let be a scheme. Then the problem is -solvable for any point set if and only if
Until the end of the subsection, we study the bivariate case again.
@@
Suppose that a scheme with non-increasing sequence of multiplicities, satisfies the condition: Then is regular if only if
Proof.
For denote
Next proposition is an extension of Proposition 1.12:
@@
Suppose is a set of points with where is an irreducible conic. Then we have that
where and
@@
The above inequality, in view of irreducibility of implies that the -condition is satisfied for the problem .
Thus, in view of Theorem (1.18), we obtain that is -solvable. Similarly we obtain that the problem is -solvable.
Now, consider the following linear spaces:
Note that Since the problems and are solvable, we get from Corollary 1.9, that
@@
Then implies that While implies that
Now, let Note that is exact. The proof of Proposition 1.21 implies that the problem is solvable, provided that with any irreducible conic. The latter condition in this case means that no three points of are collinear. Thus the problem is correct:
Note that also (see the above proof) is correct, since if then Therefore Proposition 1.21 is valid also in the case and the statement (1.13), i.e., reduces to
@@
Note that, according to Corollary 1.20, for any scheme with non-increasing sequence of multiplicities, we have that the conditions imply that is -scheme.
In the main result of this paper below we consider instead the condition which, in combination with does not imply that is -scheme. Indeed, for a counterexample, we just add a fifth multiplicity to the scheme of the above remark.
This is why, in the following theorem we have an additional condition, compared to Theorem 1.18: is -scheme.
2 The main result
@@
Then the problem is -solvable if and only if:
There is no line passing through more than points of
There is no conic passing through more than points of
@@
The direction “only if” follows from Corollary 1.15. Let us prove the direction “if”. First we consider the (main) case
Note that the above condition (ii) means that there is no conic passing through all the points of
@@
By using the condition (i) for the line passing through we get
@@
Indeed, assume that Then the only -scheme is the the Taylor scheme But the latter does not satisfy (2.1). While if then the conditions (2.1) and (2.3) imply that which is not -scheme.
@@
In the case there is only one -scheme satisfying the conditions (2.1) and (2.4). According to the condition (ii) of Theorem, the six simple points of the set do not belong to a conic. Since is exact scheme, we obtain from Proposition 1.5(ii), that is -correct.
@@
Step 2. Assume that there is a line which passes through points of i.e., the following condition is satisfied:
@@
Next, we want to apply the induction hypothesis for the problem For this let us check first that is -scheme. Indeed, by using (2.6), we get
The condition (2.1) for readily follows from the fact that passes through at least two points of
Then let us verify that the -condition is satisfied with respect to the problem i.e., there is no line containing points of Assume conversely that there is such a line Then evidently and, in view of the relation
we obtain that contains points from Since contains points both from and we conclude that the line does not intersect at a point of Therefore, the (reducible) conic passes through points of which contradicts the condition (ii) of Theorem.
Finally let us verify the -condition for the problem Assume, by way of contradiction, that there is a conic containing points of In view of (2.7) this means that passes through all the points with
Note that if passes through at least three points of then in (2.7) we would obtain and hence the -condition is satisfied. It remains to consider the case where passes through exactly two points of , say, and In this case, in view of (2.5), we have that According to (2.4) we get that Hence we have that Thus passes through all the points of which is contradiction.
@@
Now, Proposition 1.8 implies that there is a set consisting of simple points (see (1.7) and (1.8)), such that the problem
@@
Here, by making use of Remark 1.10, we can also assume that the set contains no point belonging to the line
@@
Let us check first that is exact scheme. Indeed, in view of (2.6), we have
Next, let us prove that the problem is -correct. For this, let us assume, in view of Proposition 1.5(ii), that where and prove that
@@
Then, Proposition 1.12 implies that where
Now, in view of (2.8) and Proposition 1.5(ii), we get and hence
@@
Step 3. Assume that there is a line passing through at least points of Consider the following scheme (see (1.9))
@@
Now, Proposition 1.8 implies that there is a set consisting of simple points, such that the problem
@@
Again, by making use of Remark 1.10, we can assume also that the set contains no point belonging to the line
Then denote by a set of simple points in
@@
Next, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.5(ii), that where and prove that
@@
Now, in view of (2.10) and Proposition 1.5(ii), we get and hence
@@
From now on, in view of Step 3, we may assume that there are no three collinear points in the set
Step 4. First note that the condition (ii) of Theorem, in view of Corollary 1.7, implies that Then consider a conic passing through the points The conic is irreducible, since no three points are collinear.
@@
Now, assume conversely, that the inequality holds for some Then, in view of (2.2), we have that and therefore
@@
Thus, according to Theorem 1.18, the problem is -solvable. Therefore, Proposition 1.8 implies that there is a set consisting of simple points, such that the problem
@@
By making use of Remark 1.10 we can assume that the set contains no point belonging to the conic
Next, denote by a set of simple points in
@@
Now, let us prove that the problem is -correct. For this let us assume, in view of Proposition 1.5(ii), that where and prove that
@@
Then, in view of (2.11) and Proposition 1.5(ii), we get and hence
@@
Step 5. It remains to consider the case of general dimension.
Assume that is a set of points in dimension satisfying the and -conditions with respect to the problem Then we can project it to a set of points in dimension, with the same property. Continuing in this way we will obtain a set of points in a plane, with the same property. Now, the problem is -solvable, since Theorem has already been proven in the bivariate case. From here, by using times Remark 1.10, we obtain that the original problem is -solvable too. ∎
A counterexample.
Suppose the scheme is the following Thus (2.12) takes place. Also we have that is -scheme:
Suppose the corresponding set of points is chosen as follows: The points all belong to an irredusible conic The points and belong to a line which does not contain any other point from
It is easily seen that the conditions (i) and (ii) of Theorem 2.1 are satisfied:
Let us verify that the Hermite problem is not -solvable.
Indeed, by applying Corollary 1.13 with respect to the line we get that the problem is -solvable if and only if the problem is -solvable, where and
Now, on the conic we have points of Hence, in view of Corollary 1.15(ii), the problem is not -solvable.
From Theorem 2.1 we get
Corollary 2.2.
Suppose that a -scheme with non-increasing sequence of multiplicities, satisfies the condition: Then is regular if and only if and
Proof.
Consider the following exact scheme This scheme satisfies the conditions of above corollary, with one exception: According to Corollary 1.17 the scheme is not regular.